OpenStudy (anonymous):

I HAVE my digital circuits xam tommorow.. cnany1 plz help me out wid this problem Find the complement of the expression: (A'B + CD)E' + E

OpenStudy (anonymous):

no help :(

OpenStudy (anonymous):

Simply use D Morgan Theorem (A+B')(C'+D')E'

OpenStudy (anonymous):

I think it will be:- (A'B+CD)E'+E USING DEMORGAN >>[(A'B+CD)E'+E]' >>[(A'B+CD)'+E"".E'] i.e E"=E >>{[(A"+B')(C'+D')]+0} i.e E.E'=0 >>(A+B')(C'+D')

OpenStudy (anonymous):

get the dual of the expression and then compliment each of the variable , that is the shot cut to check ur result dual expression : (A'+B)(C+D)+EE' when each variable is complimented : ((A')' + B' ) (C' +D')+1'= (A+B')(C'+D')+0= (A+B')(C'+D')